A brass plug is to be placed in a ring made of iron. At 20 ∘C, the diameter of the plug is 8.829 cmand that of the inside of the ring is 8.819 cm . They must both be brought to what common temperature in order to fit?

What if the plug were iron and the ring brass?
Answer #1

At some temperature, lbrass=liron

∴ l=l0+l0αΔT.(brass)=l0+l0αΔT.(iron)

8.829cm+(8.829cm×18.7×10-6°C-1×ΔT)
=8.819cm+(8.819cm×12.0×10-6°C-1×ΔT)

0.01+(8.829cm×18.7×10-6°C-1×ΔT)= (8.819cm×12.0×10-6°C-1×ΔT)

ΔT= -168.71°C.

Both brass and ring must be cooled by 168.71°C.

If you start at 20 °C, you must cool them to -148.71 °C